# The Pólya-Tchebotarev problem with semiclassical external fields

### v1

@@ 3.1 Critical Measures
From the general theory of quadratic differentials [24], at each pole $z_{j}$ there is neighborhood $U$ of $z_{j}$ such that every trajectory intersecting $U$ ends up at $z=z_{j}$ in at least one direction. In particular, if $\gamma\cap U\neq\emptyset$ for some maximal trajectory $\gamma$ of $\operatorname{supp}\mu$, then this arc $\gamma$ ends at $z_{j}$. On the other hand, from the explicit expression (3.23) for the density of $\mu$ and the fact that $R(s)$ has growth of order at least $4$ when $s\to z_{j}$ over the trajectories, we see that if $\gamma$ were to extend to a pole $z_{j}$, then $\mu$ would not be a finite measure. Analogously, near logarithmic singularities of $\Phi$ the density (3.23) would blow up linearly, and therefore $\mu$ would not be finite. This shows the claimed properties (2) and (3). $R$ is a rational function and therefore the possible trajectories are critical trajectories that connects zeroes and simple poles of $R$. Since there are only a finite number of them, the union is over a finite number of arcs. ∎
@@ 3.2 Critical Sets
Variations $z\mapsto z+th(z)$ induces variations of measures, which we studied in detail in Section 3.1. In a similar way, it also defines variations of sets, that is, the analogue of Definition 3.1, as we study next.
@@
We shall denote by $\mathcal{M}_{0}$ the set of signed measures defined on $\mathbb{C}$, whose positive and negative parts belong to $\mathcal{M}_{1}(\mathbb{C})$ and have finite logarithmic energy. A sequence of finite signed measures $(\sigma_{n})$ converges vaguely to a finite signed measure $\sigma$ when
@@
for every continuous function $f$ with compact support. In what follows, we also talk about weak and vague convergence of a family of measures $(\mu_{t})=(\mu_{t})_{t\in\mathbb{R}\setminus\{0\}}$ as $t\to 0$ to a measure $\mu_{0}$. By such convergence, we mean that that $\mu_{t_{n}}\to\mu_{0}$ in the weak/vague sense, for any sequence of indices $(t_{n})$ with $t_{n}\to 0$ as $n\to\infty$, and write $\mu_{t}\stackrel{{\scriptstyle t\to 0}}{{\to}}\mu_{0}$ in the weak/vague sense. The following result is well known for compactly supported signed measures, and it was extended to allow for unbounded support in [14].
Lemma 3.14 ([14, Lemma 4.3]).
@@
and to finish the lemma we show that the right-hand side can be made arbitrarily small. We first examine the second term on the right-hand side, starting with the obvious inequality
@@
because $h(x)C_{R,0}(x)$ is continuous on $D_{R-1}\supset D_{R_{0}}$ and with support on $D_{R_{0}}\supset\operatorname{supp}h$. Finally, to handle the first term in the right-hand side of (3.25) we use some auxiliary measure-theoretic results. Thanks to Proposition A.3 we can always choose $R>R_{0}+1$ for which $\mu(\partial D_{R})=0$. In particular, using that $\partial(D_{R}\times D_{R})\subset(\partial D_{R}\times\mathbb{C})\cup(\mathbb{C}\times\partial D_{R})$, we see that $\mu\times\mu(\partial(D_{R}\times D_{R}))=0$. From the vague convergence $\mu_{n}\times\mu_{n}\to\mu\times\mu$ and Proposition A.2, we conclude
@@
Thus, to finish the proof, it is sufficient to show that $\operatorname{D}_{h}\operatorname{I}^{\varphi}(\mu_{F})-\operatorname{D}_{h}\operatorname{I}^{\varphi}(\mu_{F^{t}})=o(1)$ as $t\to 0$. Assume for a moment that $\mu_{F^{t}}\to\mu_{F}$ vaguely as $t\to 0$, fact which will be proved later. The function $h\Phi^{\prime}$ belongs to $C_{c}^{2}$, so
@@
which then concludes the proof. Finally, it remains to verify that $\mu_{F^{t}}$ vaguely converges to $\mu_{F}$ as $t\to 0$. For that we will verify that $(\mu_{F}^{t})$ converges to $\mu_{F}$ weakly, and that $\mu_{F^{t}}-\mu_{F}^{t}$ converges to the null measure vaguely.
@@
for each bounded continuous function $f$. By Proposition 3.2, the difference $D_{h}\operatorname{I}^{\varphi}(\mu_{F})-D_{h}\operatorname{I}^{\varphi}(\mu_{F^{t}})$ is bounded. Thus, Equation (3.30) implies that $\operatorname{I}^{\varphi}(\mu_{F}^{t})-\operatorname{I}^{\varphi}(\mu_{F^{t}})\to 0$ as $t\to 0$. Now, by Lemma 3.14, as $t\to 0$,
@@ 4 Solving the max-min problem
Having developed the theory of critical measures, we now go back to solving the max-min problem as posed in (2.18), for a class of contours $\mathcal{T}$ for an admissible triple $(\mathcal{P}(\mathcal{C}),\Psi,\mathcal{P}(\Theta_{\infty}))$. Following [14], such solution will be established in three steps as follows. The first step is to construct a class of sets $\mathcal{F}_{M}$ which is different but related to our initial class of contours $\mathcal{T}$. The advantage of $\mathcal{F}_{M}$ will be that it is closed under the Hausdorff metric, but as a consequence it contains rather arbitrary closed sets, not only contours. We then solve a max-min problem on $\mathcal{F}_{M}$, which will be essentially a consequence of Theorem 2.6, obtaining a max-min set $F_{0}$. The second part, which is also the longest, consists of reconstructing a contour $\Gamma_{0}\in\mathcal{T}$ from the set $F_{0}$, in such a way that their equilibrium measures coincide. This construction is done with a careful dropping of unnecessary parts of it, while preserving some geometry of the set, in a procedure that relies heavily on the results on critical measures that we previously obtained.
@@ 4.1 Finding a max-min candidate
A solution $\Gamma_{0}$ of the max-min problem should not intersect parts of the complex plane where the external field is very negative, since then $\Gamma_{0}$ would have very negative energy, and thus it cannot be the maximum over $\mathcal{T}$. We exploit this in the following, presenting an equivalent max-min problem over a subfamily of $\mathcal{T}$ whose closure on the Hausdorff metric satisfies the hypotheses of Rakhmanov’s Theorem. For $M$ large enough the set
[…]
is made of three types of curves (see Figure 5):
@@
    at each pole $z_{j}$ of $\Phi$ of order $m_{j}$ there are $m_{j}$ “tear-like” curves, which are curves starting and ending at $z_{0}$;
@@
    because $\operatorname{Im}\rho_{j}\neq 0$, at each logarithmic singularity $w_{j}$ the set $\varphi^{-1}(M)$ extends to an analytic arc ending at $w_{j}$.
To verify above one writes $z-w=Re^{i\theta}$ and study what happens when $z\to 0$, where $w$ is a singularity of $\Phi$. For example, when $w=w_{k}$ is a logarithmic singularity, then
$\varphi(z)=\log R\operatorname{Re}\rho_{k}-\theta\operatorname{Im}\rho_{k}+g(z)$
where $g$ is an analytic function around $w_{k}$. Equating identity above to $-M$ and making $R\to 0$ we see that since $\operatorname{Im}\rho_{k}\neq 0$ (see (2.7)), $\theta$ oscillates between $(0,2\pi)$, giving a spiral behavior, or around a constant, making a slightly deformed line segment.
@@
Type (i) curves for $\Phi(z)=z^{5}-1$ and $M=15$.
Type (ii) curve for $\Phi(z)=\frac{(z-\mathrm{i})^{4}}{z^{2}}$ and $M=200$.
Type (iii) curves for $\Phi(z)=z-\mathrm{i}+\log(z)$ and the choices $M=3,4,5,6,7,8$.
@@
Let $S=S_{M}$ be the union of the type (i) curves. For sufficiently large $M$, the set
[…]
is a non-empty open set whose boundary consists of a union of $N$ pairwise disjoint analytic arcs. Consider the subclass
@@
$\sup_{\Gamma\in\mathcal{T}}\operatorname{I}^{\varphi}(\Gamma)=\sup_{\Gamma\in\mathcal{T}_{M}}\operatorname{I}^{\varphi}(\Gamma)\leq\sup_{F\in\mathcal{F}_{M}}\operatorname{I}^{\varphi}(F)<+\infty$
@@
The first inequality is always true, independent of $M$, simply because $\mathcal{T}_{M}\subset\mathcal{F}_{M}$. Recall that $\mathcal{C}$ is a finite set and $\varphi$ is finite at each point in $\mathcal{C}$ (see (2.10)), and that $\mathcal{Z}$ is the set of singularities of $\Phi$. For the rest of the proof, assume that $M>0$ satisfies $-2M<\sup_{\Gamma\in\mathcal{T}}\operatorname{I}^{\varphi}(\Gamma)$ and $-M<\inf\varphi(\mathcal{C})$. At this stage we are allowing this supremum to possibly be equal to +$\infty$+ ∞. To verify the claimed equality, it suffices to show that
  • There is a connected contour $\gamma\subset\Gamma$ that intersects $\Delta_{M}$ and connects a fixed point $c$ to infinity through one of the admissible sectors. Since $\varphi(c)>-M$, by the very definition of $\Delta_{M}$ we see that $\gamma$ contains a compact connected contour $\gamma^{\prime}$ satisfying
    $\gamma^{\prime}\subset\varphi^{-1}(-\infty,-M)\quad\text{and}\quad\operatorname{diam}\gamma^{\prime}\geq 8.$ (4.2)
  • There is a connected contour $\gamma\subset\Gamma$ that intersects $\Delta_{M}$ and connects two fixed points $c_{1},c_{2}$. Again, since $\varphi(c_{1}),\varphi(c_{2})>-M$, the definition of $\Delta_{M}$ ensures that $\gamma$ contains a finite contour $\gamma^{\prime}$ satisfying (4.2).
  • There is a connected contour $\gamma\subset\Gamma$ that intersects $\Delta_{M}$ and connects infinity through two admissible sectors. Once again, there is $\gamma^{\prime}\subset\gamma$ satisfying (4.2).
In either of the cases singled out, the estimate of capacity in terms of diameter $\operatorname{cap}(\gamma^{\prime})\geq\frac{\operatorname{diam}\gamma^{\prime}}{4}\geq 2$ (see [26, Theorem 5.3.2]) implies that $\operatorname{I}(\gamma^{\prime})\leq-\log 2$. Let $\omega$ be the Robin measure of $\gamma^{\prime}$, that is, the minimizer of $\operatorname{I}$ over the class $\mathcal{M}_{1}^{\varphi\equiv 0}(\gamma^{\prime})$. Then
This holds for each $\Gamma\in\mathcal{T}\setminus\mathcal{T}_{M}$, which shows (4.1) and concludes the proof of the claimed equality. We now verify the third inequality. To that end, consider a simple closed curve $\gamma$ contained in $\mathbb{C}\setminus\Delta_{M}$ and intersecting all connected components of $\partial\Delta_{M}$ exactly once. Then $\mathbb{C}\setminus\gamma$ has two connected components. By making $M>0$ sufficiently large we make sure that the finite set $\mathcal{Z}\cup\mathcal{C}$ is contained in the bounded connected component of $\mathbb{C}\setminus\gamma$.
@@
Next, to continue, we split the analysis into two separate cases, namely when each $\Gamma\in\mathcal{T}$ is unbounded, and the remaining case when we do not enforce contours $\Gamma\in\mathcal{T}$ to be unbounded. In terms of Definition 2.2, we do not enforce contours $\Gamma\in\mathcal{T}$ to be unbounded if, and only if, $\mathcal{P}_{\Psi}=\mathcal{P}_{0}(\Theta)=\emptyset$. Thus, in the former case aforementioned, each $\Gamma\in\mathcal{T}_{M}$ possesses an unbounded connected component $\widehat{\Gamma}$ that stretches to infinity in a fixed admissible sector, meaning that $\widehat{\Gamma}$ either connects a fixed point to infinity, or connects two distinct admissible sectors at infinity, see Definition 2.2–(v). The contour $\gamma$ is constructed in such a way that $\gamma\cup\partial\Delta_{M}$ separates the points in $\mathbb{C}$ to $\infty$, and also separates different admissible sectors. Because contours in $\mathcal{T}_{M}$ must not intersect $\overline{\Delta}_{M}$, it follows that $\widehat{\Gamma}\cap\gamma\neq\emptyset$. This property is preserved under taking closure with respect to the Hausdorff metric, that is, every $F\in\mathcal{F}_{M}$ possesses a connected component $\widehat{F}$ for which $\widehat{F}\cap\gamma\neq\emptyset$.
@@
$\operatorname{I}^{\varphi}(F)\leq\operatorname{I}^{\varphi}(F_{0})\leq\operatorname{I}(F_{0})+2\sup_{z\in F_{0}}\varphi(z)\leq 2\sup_{R-4\leq|z|\leq R}\varphi(z)$
@@
The last term is finite and independent of $F$. This finishes the proof. ∎
@@
$\mathcal{F}_{M}$ is compact in the Hausdorff metric, each one of its elements has at maximum $|\mathcal{P}(C)|+|\mathcal{P}_{0}(\Theta)\setminus\mathcal{P}_{\Psi}|$ connected components and the supremum is finite by Proposition 4.1, so the result follows from Theorem 2.6. ∎
Any $F_{0}\in\mathcal{F}_{M}$ who maximizes the energy functional on $\mathcal{F}_{M}$ will be called a maximizing candidate. Note that at the moment a maximizing candidate is not guaranteed to be a solution to our max-min problem of interest: convergence in the Hausdorff metric does not preserve contours, so $F_{0}$ is not necessarily an element of $\mathcal{T}$. Our next step is to ensure that maximizing candidates do not come to the “boundary” of $\mathcal{F}_{M}$, to ensure we are able to apply the variational theory developed in Section 3. Recall that the perturbation $F^{t}$ of a set $F$ was introduced in Definition 3.12.
@@
Let $F_{0}\in\mathcal{F}_{M}$ be a maximizing candidate. For any given $h\in\mathcal{H}$ there exists $t_{0}>0$ such that $F_{0}^{t}\in\mathcal{F}_{M}$, for every $t\in(-t_{0},t_{0})$. In particular, every maximizing candidate is a critical set.

### v2

@@ 3.1 Critical Measures
From the general theory of quadratic differentials [29], at each pole $z_{j}$ there is neighborhood $U$ of $z_{j}$ such that every trajectory intersecting $U$ ends up at $z=z_{j}$ in at least one direction. In particular, if $\gamma\cap U\neq\emptyset$ for some maximal trajectory $\gamma$ of $\operatorname{supp}\mu$, then this arc $\gamma$ ends at $z_{j}$. On the other hand, from the explicit expression (3.23) for the density of $\mu$ and the fact that $R(s)$ has growth of order at least $4$ when $s\to z_{j}$ over the trajectories, we see that if $\gamma$ were to extend to a pole $z_{j}$, then $\mu$ would not be a finite measure. Analogously, near $w_{i}\in\mathcal{W}$ whose residue of $\sqrt{-R(z)}$ is not purely imaginary, the density (3.23) would blow up linearly, and therefore $\mu$ would not be finite. This shows the claimed properties (1) and (2).
The quadratic differential $-R(z)\mathrm{d}z^{2}$ has double poles at the points $w_{j}$’s, and around these points, as said, the trajectories consist of closed curves winding around $w_{j}$. As such, it could be possible that the support of a critical measure contains infinitely many of these closed trajectories. However, if in addition we assume that $\mu$ is also an equilibrium measure, then we can rule out such possibility, as claimed by the next proposition. For our final goal of understanding max-min problems, the critical measures of relevance will always be equilibrium measures as well, and the next proposition will become particularly useful.
Proposition 3.12.
Assume that $\mu$ is a critical measure that is also an equilibrium measure on the external field $\varphi=\operatorname{Re}\Phi$ for a set $S$. Suppose that the support of $\mu$ contains a connected component $\Gamma$ that is a trajectory that winds around a singularity $w_{j}$. Let $U$ be the bounded connected component of $\mathbb{C}\setminus\Gamma$. Then $\operatorname{supp}\mu\cap U=\emptyset$. In particular, a critical measure that is an equilibrium measure has at most one connected component winding around a $w_{j}\in\mathcal{W}$.
Proof.
Trajectories that wind around a singularity can only be found when $w_{j}$ is a double pole of $R$ with residue of $\sqrt{-R(z)}$ purely imaginary, and in this case the whole neighborhood $U$ is composed of trajectories with the same behavior. For such a neighborhood $U$, one has
$R(s)=\frac{R_{0}(s)}{(s-w_{j})^{2}},\quad R_{0}(w_{j})\neq 0$
where $R_{0}$ is analytic on $U$, thus the function $\sqrt{R}$ is well defined and meromorphic on $U$.
For $\delta>0$ let
$U_{\delta}=U\setminus B(z_{0},\delta).$
We claim that for sufficiently small $\delta<0$, the set $\operatorname{supp}\mu\cap U_{\delta}$ is composed of a finite amount of trajectories. To prove that first note that if $\eta$ is a trajectory contained in $U_{\delta}\cap\operatorname{supp}\mu$, then
$\mu(\eta)=\int_{\eta}|\sqrt{R(s)}|\,|\mathrm{d}s|\geq 2\pi\delta\cdot\inf_{s\in\overline{U_{\delta}}}|\sqrt{R(s)}|>0.$
Therefore, if $\eta_{n}$ denotes an enumerable pair-wise disjoint family of trajectories contained in $U_{\delta}\cap\operatorname{supp}\mu$, then
$\mu(\mathbb{C})\geq\sum_{n\geq 1}\mu(\eta_{n})=+\infty$
a contradiction. This proves the claim.
We now prove the proposition. Assume that $\operatorname{supp}\mu\cap U\neq\emptyset$. Since $\operatorname{supp}\mu$ is composed of trajectories, there exist $\eta\subset U$ a trajectory that is entirely contained in $\operatorname{supp}\mu$. Let $\Gamma^{\prime}$ be the trajectory of $\operatorname{supp}\mu\cap U$ that is the closest to $\Gamma$, i.e., that there is no trajectory of $\operatorname{supp}\mu\cap U$ in the “ring-like” domain whose boundary is given by $\Gamma\cap\Gamma^{\prime}$.
For each $y\in\Gamma^{\prime}$, let $\alpha=\alpha_{y}$ be the orthogonal trajectory passing through $y$. We see $\alpha$ as a oriented contour that intersects $\Gamma$ once at a point $x$, passes through $y$ and then goes to $w_{j}$.
The function
$f(z)=\int_{x}^{z}\sqrt{R(s)}\,\mathrm{d}s.$
is well defined in the “ring-like” domain whose boundary is given by $\Gamma\cup\Gamma^{\prime}$. The identity
$\operatorname{Re}f(z)=U^{\mu}(z)+\varphi(z)-l_{x},\quad l_{x}=U^{\mu}(x)+\varphi(x).$
immediately follows from the definition of $f$ and the algebraic equation satisfied by $C^{\mu}$, $\varphi$ and $R$. Since $\operatorname{Im}f$ is constant along $\alpha$, $\operatorname{Re}f$ is monotone along $\alpha$ by the Cauchy-Riemann equations, thus, for all $z\in\alpha$
$0=\operatorname{Re}f(x)<\pm\operatorname{Re}f(z)=\pm\left[U^{\mu}(z)+\varphi(z)-l_{x}\right]$
where the sign depends on the monotonicity of $\operatorname{Re}f$. In any case, identity above contradicts the Euler-Lagrange variational conditions (2.4) for $\mu$ when one takes $z=y$. Thus $\operatorname{supp}\mu\cap U=\emptyset.$
Corollary 3.13.
Let $\mu$ be a critical equilibrium measure. The support of $\mu$ is a compact set made of a finite union of analytic arcs.
Proof.
In Corollary 3.11 it was shown that the support of $\mu$ is compact, made of a union of analytic curves, composed of trajectories of the quadratic differential $-R(z)\mathrm{d}z^{2}$, not containing any pole of $\Phi$, nor double poles of $R(z)$ whose residue of $\sqrt{-R(z)}$ is not purely imaginary.
Since $R$ is a rational function, the remaining possible trajectories are critical trajectories that connect zeroes and simple poles of $R$, and trajectories that wind around a $w_{j}$. By the Proposition 3.12, the amount of curves of the later case is bounded by $|\mathcal{W}|$. ∎
@@ 3.2 Critical Sets
Variations $z\mapsto z+th(z)$ induce variations of measures, which we studied in detail in Section 3.1. In a similar way, it also defines variations of sets, that is, the analogue of Definition 3.1, as we study next.
@@
We shall denote by $\mathcal{M}_{0}$ the set of signed measures defined on $\mathbb{C}$, whose positive and negative parts belong to $\mathcal{M}_{1}(\mathbb{C})$ and have finite logarithmic energy. A sequence of finite signed measures $(\sigma_{n})$ converges vaguely (resp. weakly) to a finite signed measure $\sigma$ when
@@
for every compactly supported continuous function $f$ (resp. for every bounded continuous function $f$).
In what follows, we also talk about weak and vague convergence of a family of measures $(\mu_{t})=(\mu_{t})_{t\in\mathbb{R}\setminus\{0\}}$ as $t\to 0$ to a measure $\mu_{0}$. By such convergence, we mean that that $\mu_{t_{n}}\to\mu_{0}$ in the weak/vague sense, for any sequence of indices $(t_{n})$ with $t_{n}\to 0$ as $n\to\infty$, and write $\mu_{t}\stackrel{{\scriptstyle t\to 0}}{{\to}}\mu_{0}$ in the weak/vague sense.
The following result is well known for compactly supported signed measures, and it was extended to allow for unbounded support in [18].
Lemma 3.16 ([18, Lemma 4.3]).
@@
and to finish the lemma we show that the right-hand side can be made arbitrarily small.
We first examine the second term on the right-hand side, starting with the obvious inequality
@@
because $h(x)C_{R,0}(x)$ is continuous on $D_{R-1}\supset D_{R_{0}}$ and with support on $D_{R_{0}}\supset\operatorname{supp}h$.
Finally, to handle the first term in the right-hand side of (3.25) we use some auxiliary measure-theoretic results. Thanks to Proposition A.3 we can always choose $R>R_{0}+1$ for which $\mu(\partial D_{R})=0$. In particular, using that $\partial(D_{R}\times D_{R})\subset(\partial D_{R}\times\mathbb{C})\cup(\mathbb{C}\times\partial D_{R})$, we see that $\mu\times\mu(\partial(D_{R}\times D_{R}))=0$. From the vague convergence $\mu_{n}\times\mu_{n}\to\mu\times\mu$ and Proposition A.2, we conclude
@@
Thus, to finish the proof, it is sufficient to show that $\operatorname{D}_{h}\operatorname{I}^{\varphi}(\mu_{F})-\operatorname{D}_{h}\operatorname{I}^{\varphi}(\mu_{F^{t}})=o(1)$ as $t\to 0$.
Assume for a moment that $\mu_{F^{t}}\to\mu_{F}$ vaguely as $t\to 0$, fact which will be proved later. The function $h\Phi^{\prime}$ belongs to $C_{c}^{2}$, so
@@
which then concludes the proof.
Finally, it remains to verify that $\mu_{F^{t}}$ vaguely converges to $\mu_{F}$ as $t\to 0$. For that we will verify that $(\mu_{F}^{t})$ converges to $\mu_{F}$ weakly, and that $\mu_{F^{t}}-\mu_{F}^{t}$ converges to the null measure vaguely.
@@
for each bounded continuous function $f$.
By Proposition 3.2, the difference $D_{h}\operatorname{I}^{\varphi}(\mu_{F})-D_{h}\operatorname{I}^{\varphi}(\mu_{F^{t}})$ is bounded. Thus, Equation (3.30) implies that $\operatorname{I}^{\varphi}(\mu_{F}^{t})-\operatorname{I}^{\varphi}(\mu_{F^{t}})\to 0$ as $t\to 0$. Now, by Lemma 3.16, as $t\to 0$,
@@ 4 Solving the max-min problem
Having developed the theory of critical measures, we now go back to solving the max-min problem as posed in (2.18), for a class of contours $\mathcal{T}$ for an admissible triple $(\mathcal{P}(\mathcal{C}),\Psi,\mathcal{P}(\Theta_{\infty}))$. Following [18], such solution will be established in three steps as follows.
The first step is to construct a class of sets $\mathcal{F}_{M}$ which is different but related to our initial class of contours $\mathcal{T}$. The advantage of $\mathcal{F}_{M}$ will be that it is closed under the Hausdorff metric, but as a consequence it contains rather arbitrary closed sets, not only contours. We then solve a max-min problem on $\mathcal{F}_{M}$, which will be essentially a consequence of Theorem 2.6, obtaining a max-min set $F_{0}$.
The second part, which is also the longest, consists of reconstructing a contour $\Gamma_{0}\in\mathcal{T}$ from the set $F_{0}$, in such a way that their equilibrium measures coincide. This construction is done with a careful dropping of unnecessary parts of it, while preserving some geometry of the set, in a procedure that relies heavily on the results on critical measures that we previously obtained.
@@ 4.1 Finding a max-min candidate
A solution $\Gamma_{0}$ of the max-min problem should not intersect parts of the complex plane where the external field is very negative, since then $\Gamma_{0}$ would have very negative energy, and thus it cannot be the maximum over $\mathcal{T}$. We exploit this in the following, presenting an equivalent max-min problem over a subfamily of $\mathcal{T}$ whose closure on the Hausdorff metric satisfies the hypotheses of Rakhmanov’s Theorem.
For $M$ large enough the set $\varphi^{-1}(-M)$ is made of three types of curves (see Figure 5):
@@
    at each pole $z_{j}$ of $\Phi$ of order $m_{j}$ there are $m_{j}$ “tear-like” curves, which are curves starting and ending at $z_{j}$;
@@
    If $\rho_{k}>0$, there is a circle-like curve winding around $w_{k}$.
To verify above one writes $z-w=Re^{i\theta}$ and study what happens when $z\to w$, where $w$ is a singularity of $\Phi$. For example, when $w=w_{k}$, then
$\varphi(z)=\rho_{k}\log R+g(z)$
where $g$ is analytic at $w_{k}$. Equating identity above to $-M$ and making $R\to 0$ we see that if $\rho_{k}>0$, the solution set is a curve winding around $w_{k}$, while if $\rho_{k}<0$, there is a neighborhood of $w_{k}$ in which the solution set is empty.
@@
Type (iii) curves for $\Phi(z)=z-1+2\log(z)$ and the choices $M=3,4,5,6,7,8$.
@@
Let $S=S_{M}$ be the union of the type (i) curves. Define
$r_{M}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\inf\{|z|;z\in S\}.$
For sufficiently large $M$, the set
$\Delta_{M}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\varphi^{-1}(-\infty,-M)\cap\{z\in\mathbb{C};|z|>r_{M}\text{ and }\operatorname{dist}(z,S)>8\}$
is a non-empty open set whose boundary consists of a union of $N$ pairwise disjoint analytic arcs.
Consider the subclass
@@
with respect to the Hausdorff metric on closed subsets of $\mathbb{C}$ with the chordal distance.
@@
$\sup_{\Gamma\in\mathcal{T}}\operatorname{I}^{\varphi}(\Gamma)=\sup_{\Gamma\in\mathcal{T}_{M}}\operatorname{I}^{\varphi}(\Gamma)\leq\sup_{F\in\mathcal{F}_{M}}\operatorname{I}^{\varphi}(F)<+\infty$ (4.1)
@@
The first inequality is always true, independent of $M$, simply because $\mathcal{T}_{M}\subset\mathcal{F}_{M}$.
We start now verifying the second inequality. Recall that $\mathcal{C}$ is a finite set and $\varphi$ is finite at each point in $\mathcal{C}$ (see (2.10)), and that $\mathcal{Z}$ is the set of singularities of $\Phi$. For the rest of the proof, assume that $M>0$ satisfies $-2M<\sup_{\Gamma\in\mathcal{T}}\operatorname{I}^{\varphi}(\Gamma)$ and $-M<\inf\varphi(\mathcal{C})$. At this stage we are allowing this supremum to possibly be equal to +$\infty$+ ∞.
Consider a simple closed curve $\gamma$ contained in $\mathbb{C}\setminus\Delta_{M}$ and intersecting all connected components of $\partial\Delta_{M}$ exactly once. Then $\mathbb{C}\setminus\gamma$ has two connected components. By making $M>0$ sufficiently large we make sure that the finite set $\mathcal{Z}\cup\mathcal{C}$ is contained in the bounded connected component of $\mathbb{C}\setminus\gamma$.
@@
Next, to continue, we split the analysis into two separate cases, namely when each $\Gamma\in\mathcal{T}$ is unbounded, and the remaining case when we do not enforce contours $\Gamma\in\mathcal{T}$ to be unbounded. In terms of Definition 2.2, we do not enforce contours $\Gamma\in\mathcal{T}$ to be unbounded if, and only if, $\mathcal{P}_{\Psi}=\mathcal{P}_{0}(\Theta)=\emptyset$.
Thus, in the former case aforementioned, each $\Gamma\in\mathcal{T}_{M}$ possesses an unbounded connected component $\widehat{\Gamma}$ that stretches to infinity in a fixed admissible sector, meaning that $\widehat{\Gamma}$ either connects a fixed point to infinity, or connects two distinct admissible sectors at infinity, see Definition 2.2–(v). The contour $\gamma$ is constructed in such a way that $\gamma\cup\partial\Delta_{M}$ separates $\infty$ from the points of the bounded connected component of $\mathbb{C}\setminus\gamma$, and also separates different admissible sectors from each other. Because contours in $\mathcal{T}_{M}$ must not intersect $\overline{\Delta}_{M}$, it follows that $\widehat{\Gamma}\cap\gamma\neq\emptyset$. This property is preserved under taking closure with respect to the Hausdorff metric, that is, every $F\in\mathcal{F}_{M}$ possesses a connected component $\widehat{F}$ for which $\widehat{F}\cap\gamma\neq\emptyset$.
@@
$\operatorname{I}^{\varphi}(F)\leq\operatorname{I}^{\varphi}(F_{0})\leq\operatorname{I}(F_{0})+2\sup_{z\in F_{0}}\varphi(z)\leq 2\sup_{R-4\leq|z|\leq R}\varphi(z)$ (4.2)
The last term is finite and independent of $F\in\mathcal{F}_{M}$, thus the inequality $\sup_{F\in\mathcal{F}_{M}}\operatorname{I}^{\varphi}(F)<+\infty$ follows.
Now assume that contours in $\mathcal{T}$ do not necessarily extend to $\infty$, this implies that there exists $\mathcal{P}_{0}\subset\mathcal{P}(\mathcal{C})$ with $|\mathcal{P}_{0}|\geq 2$. We fix distinct points $c,\hat{c}\in\mathcal{P}_{0}$, and choose $r>0$ such that $\overline{D_{8r}(c)}\cap(\mathcal{Z}\cup\mathcal{C})=\{c\}$.
@@
The last term is finite and independent of $F$.
If there is no uniformly bounded sequence $\Gamma_{n}$ with Hausdorff limit $F$, there exists a sequence $\Gamma_{n}$ converging to $F$ in the Hausdorff metric such that
$\Gamma_{n}\cap\partial D_{R}\neq\emptyset,\quad\Gamma_{n}\cap\partial D_{R-4}\neq\emptyset$
and we can use the same arguments used to obtain the bound given by (4.2). This finishes the proof of the lemma.
We now prove the equality in (4.1), it suffices to show that
  • There is a connected contour $\gamma\subset\Gamma$ that intersects $\overline{\Delta_{M}}$ and connects a fixed point $c$ to infinity through one of the admissible sectors.
  • There is a connected contour $\gamma\subset\Gamma$ that intersects $\overline{\Delta_{M}}$ and connects two fixed points $c_{i}$, $c_{j}$.
  • There is a connected contour $\gamma\subset\Gamma$ that intersects $\overline{\Delta_{M}}$ and connects infinity through two admissible sectors.
In each of the cases there exists a point $x\in\gamma$ such that $\varphi(x)>-M$. By following a contour starting at $x$ that intersects $\overline{\Delta_{M}}$, we obtain a subcontour $\gamma^{\prime}$ such that
$\gamma^{\prime}\cap S\neq\emptyset,\quad\gamma^{\prime}\subset\varphi^{-1}(-\infty,-M)\quad\text{and}\quad\gamma^{\prime}\cap\overline{\Delta_{M}}\neq\emptyset.$ (4.4)
In particular, $\operatorname{diam}\gamma^{\prime}>8$, since $\operatorname{dist}(S,\overline{\Delta_{M}})\geq 8$. The estimate of capacity in terms of diameter $\operatorname{cap}(\gamma^{\prime})\geq\frac{\operatorname{diam}\gamma^{\prime}}{4}\geq 2$ (see [31, Theorem 5.3.2]) implies that $\operatorname{I}(\gamma^{\prime})\leq-\log 2$. Let $\omega$ be the Robin measure of $\gamma^{\prime}$, that is, the minimizer of $\operatorname{I}$ over the class $\mathcal{M}_{1}^{\varphi\equiv 0}(\gamma^{\prime})$. Then
This holds for each $\Gamma\in\mathcal{T}\setminus\mathcal{T}_{M}$, which shows (4.3) and concludes the proof of the claimed equality.
@@
$\mathcal{F}_{M}$ is compact in the Hausdorff metric, each one of its elements has at maximum $|\mathcal{P}(C)|+|\mathcal{P}_{0}(\Theta)\setminus\mathcal{P}_{\Psi}|$ connected components and the supremum is finite by Proposition 4.1, so the result follows from Theorem 2.6.
Any $F_{0}\in\mathcal{F}_{M}$ who maximizes the energy functional on $\mathcal{F}_{M}$ will be called a maximizing candidate. Note that at the moment a maximizing candidate is not guaranteed to be a solution to our max-min problem of interest: convergence in the Hausdorff metric does not preserve contours, so $F_{0}$ is not necessarily an element of $\mathcal{T}$.
Our next step is to ensure that maximizing candidates do not come to the “boundary” of $\mathcal{F}_{M}$, to ensure we are able to apply the variational theory developed in Section 3. Recall that the perturbation $F^{t}$ of a set $F$ was introduced in Definition 3.14.
@@
Let $F_{0}\in\mathcal{F}_{M}$ be a maximizing candidate. For any given $h\in\mathcal{H}$ there exists $t_{0}>0$ such that $F_{0}^{t}\in\mathcal{F}_{M}$, for every $t\in(-t_{0},t_{0})$. In particular, every maximizing candidate is a critical set, whose equilibrium measure is compactly supported, and composed of finitely many analytic arcs.